There are 4 boys and 4 girls are to be sit in a line such that boys and girls are alternate. Find how many ways ?

 

Grade:12

1 Answers

Arun
6 years ago
Dear Harish
 
The number of ways are 5C4* 4C4* 4! = 5*24 = 120 ways of arranging the 4 boys and 4 girls so that they are in alternate order
 
Regards
Arun (askIITians forum expert)